40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jun 20 00:59:02 2011, kiwamu, Summary, General, Weekly report 
    Reply  Mon Jun 27 10:04:29 2011, kiwamu, Summary, General, Weekly report 
       Reply  Mon Jul 4 14:27:35 2011, kiwamu, Summary, General, Weekly report 
          Reply  Mon Jul 25 11:02:19 2011, kiwamu, Summary, General, Weekly report 
             Reply  Mon Jul 25 11:04:22 2011, kiwamu, Summary, General, Weekly report 
                Reply  Mon Jul 25 11:06:38 2011, kiwamu, Summary, General, Weekly report 
                   Reply  Mon Aug 1 22:30:59 2011, kiwamu, Summary, General, Weekly report 
Message ID: 5026     Entry time: Mon Jul 25 11:02:19 2011     In reply to: 4936     Reply to this: 5027
Author: kiwamu 
Type: Summary 
Category: General 
Subject: Weekly report 

 Summary of the week ending July 10th.  Number of elog entries = 21

- SUS

 + The cutoff frequency of the high pass filters for the damping were set to 30Hz.
 + Turned off all the BounceRoll filters.
 + The BS oplev was checked and seemed healthy.
 

- LSC


 + All the measred data of the LSC whitening filters were fit.
 + All the zpk parameters are recorded on the wiki.

- ABSL

 

+ The setup completed.  
 + The freqeucy-lock of the ABSL laser was achieved with UGF of ~ 40kHz.
 + The temperature of the ABSL laser was adjusted to be 47.25 deg
 

- ALS

 (Fiber experiment)
 + The I-P curve of the ETMY laser was measred.
 + The current set point is 1.8 [A], which used to be 1.5 [A], corresponding to the output of power of 197 [mW] and 390 [mW] respectively.
 

ELOG V3.1.3-